grace is four times as old as vicky. five yrs ago, vicky was thirty yrs younger than grace. find their ages. Answer by mananth(16949) 
 let vicky's age =x
grace = 4x
..
5 years ago their ages were
vicky x-5
grace 4x-5
..
4x-5 -30= x-5
4x-35= x-5
3x=30
x=10 vicky"s age now
Grace's age = 40